OpenStudy (anonymous):

Use the properties of logarithms to simplify Log(4) (4x^2), where x>0. The answer I got was Log(4) (x^2), is this correct?

OpenStudy (solomonzelman):

you are thinking correctly, but you made a little error. \(\normalsize\color{blue}{ Log_4(4x^2)= Log_44+Log_4x^2= 1+Log_4x^2 =1+2Log_4x}\)

OpenStudy (anonymous):

Thank you so much for the clarification.

OpenStudy (solomonzelman):

Anytime !